For the circuit show below, use superposition to solve for V 6 ΜΩ 40 μΑ 1 ΜΩ + V - 75 μA (1) X4 MΩ 200 kΩ
Answer #1

given circuit 404A ime 6m2 L.WMT- +V- LWW 200ķh 754A ► 24 m2 E3v Find V By Superposition theorem Superposition theorem :- InNOTE while Applying superposition theorem each Sub circuit as only one independent source the other independent sources De enpor: Considered 3 volts only Im2 6m2 w tv WW 2008-th . white Z - 3y whug 1. IM 2 _ MWWW kook-2 ty- 3V z whit iMh WWW ty 10.9mBy voltage Division Rule: reoltage Across ime is im. =-3x - : 102M-2 tim.ch =-3x II 12myce -3 V = -0.267 volteStep-2 considered 75.44 only imar Gyer Two w MM 75.46 2006. 4 m2 ime 6m. 200*2ty- 54 me 54A 26-2me 15MA Q 4 + L V imarBy cuorent Division Rule. Te = 7540 x- 4me 4m2 +6-2..2 time = TEMA X 4mot. 11.2 más I = 15x4 HA 1102 T = 26.78 el A i voltageStep-3 considered 404A only Thi thoto whug Mooit v 200K who olur. throh + A - how to 2005-2 whil TV - 4m2 YouA O E 200k 3 omeу 1, 3, мя ЗООК- фоцит lome By curgent Division Rule Т. = 40 съК- x 2 2005-2 +1 М 2 том. - 4-оя хо» м-п- - О» 2 м-, + Јmn - Tie V= I xime 6 = 0.71X10 X10 V = 0.71 volts 6 By Superposition theorem Voltage heroes ime is sum) All three cases - voltages